OpenStudy (anonymous):

An aircraft makes a two way trip to a distance x miles away. The way there it travels at 240 m/hour, the way back 320 m/h. It was away for 35 min. What is x?

OpenStudy (anonymous):

v=d/t t=vd t1=going out t2=coming back t1=240mph*d t2=320mph*d Total=35min =0.58h 0.58=240d+320d Solve for d and that is the distance traveled.

OpenStudy (anonymous):

what is v?

OpenStudy (anonymous):

velocity. speed. distance traveled divided by time.

OpenStudy (anonymous):

Thankyou!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if v = d/t then t = d/v

OpenStudy (anonymous):

Well, hell! You're absolutely correct! You need to correct that and then it should work out. Good eye!

OpenStudy (anonymous):

so would it be: t1=d/240mph t2=d/320mph 0.58=d/240+d/320 and then solve for d to get the number of hours?

OpenStudy (anonymous):

Yes.